### v2

# On the small boundary property, $\mathcal{Z}$-absorption, and Bauer simplexes

Abstract.
Let $X$ be a metrizable compact space, and let $\Delta$ be a closed set of Borel probability measures on $X$. We study the small boundary property of the pair $(X,\Delta)$. In particular, it is shown that $(X,\Delta)$ has the small boundary property if it has a certain divisibility property.
As an application, it is shown that, if $A$ is the crossed product C*-algebra $\mathrm{C}(X)\rtimes\mathbb{Z}^{d}$, where $(X,\mathbb{Z}^{d})$ is a free minimal topological dynamical system, or if $A$ is an AH algebra with diagonal maps, then $A$ is $\mathcal{Z}$-absorbing if the set of extreme tracial states is compact, regardless of its dimension.
Key words and phrases:
Classification of C*-Algebras, Small Boundary Property
1991 Mathematics Subject Classification:
46L35, 37B02
1. Introduction
The small boundary property was introduced in [11] as a dynamical systems analogue of the usual definition of zero dimensional space. It was shown to be equivalent to zero mean dimension ([10]; see [6] for $\mathbb{Z}^{d}$-actions), and implies the $\mathcal{Z}$-stability of the crossed product C*-algebra $\mathrm{C}(X)\rtimes\mathbb{Z}$ ([5]). In this paper, we shall formulate the small boundary property for a pair $(X,\Delta)$ (see Definition 2.2), where $X$ is a compact metrizable space and $\Delta$ is a closed set of Borel probability measures on $X$. This clearly includes the dynamical system case where $\Delta$ consists of the invariant probability measures, but also includes the examples where $\mathrm{C}(X)$ is a Cartan subalgebra of $A$ and $\Delta$ is the trace simplex of $A$.
As suggested by the arguments of [10] and [6], the small boundary property can be characterized in the context of the uniform trace norm:
Theorem 1.1 (Theorem 2.9).
$(X,\Delta)$ has the (SBP) if, and only if, for any continuous real-valued function $f:X\to\mathbb{R}$ and any $\varepsilon>0$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{2,\Delta}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$.
This property of $\mathrm{C}(X)$ (with respect to $\Delta$) is similar to the property of real rank zero of $\mathrm{C}(X)$, which, for an arbitrary C*-algebra, asserts that any self-adjoint element can be approximated by invertible self-adjoint elements. Indeed, the (SBP) can be characterized by zero real rank of the sequence algebra of $\mathrm{C}(X)$ modulo the trace kernel of $\Delta$ (Theorem 2.12; see also Theorem 2.16).
In order to apply this theorem, consider the following property, reminiscent of uniform property $\Gamma$ (Definition 2.1 of [2]).
Definition 1.2 (Definition 3.2).
The pair $(X,\Delta)$ will be said to be approximately divisible if there is $K>0$ such that for each $n\in\mathbb{N}$, there is a partition of unity (mutually orthogonal projections with sum $1$)
$p_{1},p_{2},...,p_{n}\in\ell^{\infty}(D)/J_{2,\omega,\Delta}$
such that
$\tau(p_{i}ap_{i})\leq\frac{1}{n}K\tau(a),\quad a\in D^{+},\ \tau\in\Delta_{\omega},$
where $D=\mathrm{C}(X)$.
This property turns out to imply the (SBP):
Theorem 1.3 (Theorem 3.5).
If $(X,\Delta)$ is approximately divisible, then $(X,\Delta)$ has the (SBP).
In the case that $\mathrm{C}(X)$ is the canonical subalgebra of $A=\mathrm{C}(X)\rtimes\mathbb{Z}^{d}$, where $\mathbb{Z}^{d}$ acts minimally and freely on $X$, or in the case that $\mathrm{C}(X)$ is the diagonal subalgebra of $A$, where $A$ is an AH algebra with diagonal maps, if $\Delta=\mathrm{T}(A)$ is a Bauer simplex (i.e., $\partial\Delta$ is compact), then $(X,\Delta)$ is approximately divisible (see Lemma 4.5 below). Therefore by Theorem 1.3, it must have the (SBP). As we shall show, the C*-algebra $A$ in these cases must be $\mathcal{Z}$-absorbing (Theorem 4.6 and Proposition 4.10). Compared to the literature where Bauer simplexes are used in the classification of C*-algebras (for instance, [8], [16], and [15]), we do not use W*-bundles, and we make no assumption on the topological dimension of $\partial\Delta$.
Theorem 1.4 (cf. Theorems 4.6 and 4.9).
Let $A$ be a C*-algebra as above. If $\mathrm{T}(A)$ is a Bauer simplex, then $(X,\Delta)$ has the (SBP) and $A\cong A\otimes\mathcal{Z}$, where $\mathcal{Z}$ is the Jiang-Su algebra.
Acknowledgements
The results of this note were reported in the Operator Algebra Seminar at the Fields Institute in July, 2023 ([1]). The authors thank the Fields Institute for support. The research of the first named author was supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant, and the research of the second named author was supported by a U.S. National Science Foundation grant (DMS-1800882) and a Simons Foundation grant (MP-TSM-00002606).
2. Small boundary property
2.1. Small boundary property
Lemma 2.1.
Let $X$ be a compact metrizable space, and let $\Delta$ be a set of Borel probability measures. The following properties are equivalent:
  1. (1)
    for any subsets $V\subseteq U$, where $V$ is closed and $U$ is open, there is a closed set $V^{\prime}$ such that $V\subseteq V^{\prime}\subseteq U$, and $\mu(\partial V^{\prime})=0$, for all $\mu\in\Delta$.
  2. (2)
    for any subsets $V\subseteq U$, where $V$ is closed and $U$ is open, there is an open set $V^{\prime}$ such that $V\subseteq V^{\prime}\subseteq U$, and $\mu(\partial V^{\prime})=0$, for all $\mu\in\Delta$.
  3. (3)
    for any $x\in X$ and any open set $U\ni x$, there is a closed neighbourhood $V\subseteq U$ of $x$ such that $\mu(\partial V)=0$, for all $\mu\in\Delta$.
  4. (4)
    for any $x\in X$ and any open set $U\ni x$, there is an open neighbourhood $V\subseteq U$ of $x$ such that $\mu(\partial V)=0$, for all $\mu\in\Delta$.
Proof.
Since $\overline{\mathrm{int}(W)}\setminus\mathrm{int}(W)\subseteq\partial W=W\setminus\mathrm{int}(W)$ for any closed set $W$, (1$\Rightarrow$(2) and (3$\Rightarrow$(4) are immediate.
For (2$\Rightarrow$(3), choose an open set $U^{\prime}$ with $\overline{U^{\prime}}\subseteq U$. Since $\{x\}$ is closed, by $(2)$, there is an open set $V\ni x$ such that $V\subseteq U^{\prime}$ and $\mu(\partial V)=0$ for all $\mu\in\Delta$. Then the closed neighbourhood $\overline{V}$ of $x$ satisfies (3).
For (4$\Rightarrow$(1), choose an open set $U^{\prime}$ with $V\subseteq U^{\prime}$ and $\overline{U^{\prime}}\subseteq U$. Since $V$ is compact, by (4), there is an open cover $\{V_{1},V_{2},...,V_{n}\}$ of $V$ such that $V_{i}\subseteq U^{\prime}$, $i=1,...,n$ and $\mu(\partial V_{i})=0$, $i=1,...,n$, and $\mu\in\Delta$. Then $V^{\prime}:=\overline{V_{1}}\cup\cdots\cup\overline{V_{n}}$ satisfies (1). ∎
Definition 2.2.
Let $X$ be a metrizable compact space, fixed for the rest of the paper, and let $\Delta$ be a closed set of Borel probability measures on $X$. Then the pair $(X,\Delta)$ (or the pair $(\mathrm{C}(X),\Delta)$, where $\Delta$ is regarded as as closed set of tracial state of $\mathrm{C}(X)$) is said to have the Small Boundary Property (SBP) if the equivalent conditions of Lemma 2.1 are satisfied.
Lemma 2.3.
Let $\Delta$ be a closed set of Borel probability measures on $X$. Let $c\geq 0$, and let $V\subseteq X$ be a closed set with $\mu(V)<c$ for all $\mu\in\Delta$. Then there is an open set $U\supseteq V$ such that $\mu(U)<c$ for all $\mu\in\Delta$.
Proof.
Assume the statement were not true. Then there would exist a decreasing sequence of open sets $(U_{i})$ with $\overline{U_{i+1}}\subseteq U_{i}$ and $\bigcap_{i=1}^{\infty}U_{i}=V$ and a sequence $(\mu_{i})$ in $\Delta$ such that $\mu_{i}(U_{i})\geq c$ for all $i=1,2,...$ (and hence $\mu_{k}(U_{i})\geq c$ for all $k\geq i$, $k$, $i=1,2,...$). Since $\Delta$ is closed (and so compact), passing to a subsequence, one may assume that $(\mu_{i})$ converges to a measure $\mu_{\infty}\in\Delta$. Then, since
$\mu_{\infty}(U_{i})\geq\mu_{\infty}(\overline{U_{i+1}})\geq\limsup_{k\to\infty}\mu_{k}(\overline{U_{i+1}})\geq\limsup_{k\to\infty}\mu_{k}(U_{i+1})\geq c,$
and $\mu_{\infty}\in\Delta$, one has
$c>\mu_{\infty}(V)=\mu_{\infty}(\bigcap_{i=1}^{\infty}U_{i})=\lim_{i\to\infty}\mu_{\infty}(U_{i})\geq c,$
which is absurd. ∎
The following lemma is Proposition 5.3 of [11], which was stated for dynamical systems in terms of orbit capacity. For the reader’s convenience, we include the proof here.
Lemma 2.4 (Proposition 5.3 of [11]).
If $(X,\Delta)$ has the (SBP), then, for every open cover $\alpha$ of $X$ and every $\varepsilon>0$, there is a subordinate partition of unity $\phi_{i}:X\to[0,1]$, $i=1,2,...,\left|\alpha\right|$, such that
  1. (1)
    $\sum_{i=1}^{\left|\alpha\right|}\phi_{i}(x)=1$, $x\in X$,
  2. (2)
    $\mathrm{supp}(\phi_{i})\subseteq U$ for some $U\in\alpha$, and $i=1,2,...,\left|\alpha\right|$, and
  3. (3)
    $\mu(\bigcup_{i=1}^{\left|\alpha\right|}\phi_{i}^{-1}(0,1))<\varepsilon$ for all $\mu\in\Delta$.
Proof.
List $\alpha=\{U_{1},U_{2},...,U_{\left|\alpha\right|}\}$. Using the (SBP), for each $i=1,...,\left|\alpha\right|$, one finds an open set $U^{\prime}_{i}\subseteq U_{i}$ such that $U^{\prime}_{i}$, $i=1,2,...,\left|\alpha\right|$, still form an open cover of $X$, and $\mu(\partial U^{\prime}_{i})=0$ for all $\mu\in\Delta$. By Lemma 2.3, there is $\delta>0$ such that
$\mu((\partial U^{\prime}_{i})_{\delta})<\varepsilon/\left|\alpha\right|,$
where $(\partial U^{\prime}_{i})_{\delta}$ is the $\delta$-neighbourhood of $\partial U^{\prime}$. One may assume that $\delta$ is small enough that $(\partial U^{\prime}_{i})_{\delta}\subseteq U_{i}$ for each $i=1,...,\left|\alpha\right|$.
Define
$\psi_{i}=\left\{\begin{array}[]{ll}1,&x\in U^{\prime}_{i},\\ \max(0,1-\delta^{-1}\mathrm{dist}(x,\partial U^{\prime}_{i})),&\textrm{otherwise}.\end{array}\right.$
Then, define
$\left\{\begin{array}[]{rcl}\phi_{1}(x)&=&\psi_{1}(x),\\ \phi_{2}(x)&=&\min(\psi_{2},1-\phi_{1}(x)),\\ \phi_{3}(x)&=&\min(\psi_{3},1-\phi_{1}(x)-\phi_{2}(x)),\\ &\vdots&\\ \phi_{\left|\alpha\right|}(x)&=&\min(\psi_{\left|\alpha\right|},1-\phi_{1}-\cdots-\phi_{\left|\alpha\right|-1}(x)).\end{array}\right.$
Then, $\mathrm{supp}(\phi_{i})\subseteq U_{i}$, $i=1,...,\left|\alpha\right|$, and
$\bigcup_{i=1}^{\left|\alpha\right|}\phi_{i}^{-1}(0,1)\subseteq\bigcup_{i=1}^{\left|\alpha\right|}\psi_{i}^{-1}(0,1).$
The following lemma is contained in the proof of Theorem 6.2 of [10] and also in the proof of Corollary 5.4 of [6]. We state and prove it for the convenience of the reader.
Lemma 2.5 (Section 6 of [10]; Section 5 of [6]).
$(X,\Delta)$ has the (SBP) if, and only if, for any continuous real-valued function $f:X\to\mathbb{R}$ and any $\varepsilon>0$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{\infty}<\varepsilon$,
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$.
Proof.
Assume the properties of the lemma. In order to show that $(X,\Delta)$ has the (SBP), it is enough to show that the set
$\{f\in\mathrm{C}_{\mathbb{R}}(X):\mu(f^{-1}(0))=0,\ \mu\in\Delta\}$
is dense in $\mathrm{C}_{\mathbb{R}}(X)$ (it is always a $G_{\delta}$ set; see Lemma 6.4 of [10]). Indeed, let $x\in X$ and let $U$ be an open neighbourhood of $x$, and pick a continuous function $f:X\to[-1,1]$ such that $f(x)=-1$ and $f(X\setminus U)=\{1\}$. Since the set above is dense, there is $g$ such that $\left\|g-f\right\|<1/4$ and $\mu(g^{-1}(0))=0$ for all $\mu\in\Delta$. Consider the set $V=g^{-1}([-1,0])\subseteq U$. It is a closed neighbourhood of $x$ with $\partial V=g^{-1}(0)$, and hence $\mu(\partial V)=0$ for all $\mu\in\Delta$. So $(X,\Delta)$ has the (SBP).
For each $c>0$, consider the set
$Z_{c}:=\{f\in\mathrm{C}_{\mathbb{R}}(X):\mu(f^{-1}(0))<c,\ \mu\in\Delta\}.$
Note that the properties of the lemma imply that $Z_{c}$ is dense in $\mathrm{C}_{\mathbb{R}}(X)$.
Let us show that $Z_{c}$ is also open in $\mathrm{C}_{\mathbb{R}}(X)$. Indeed, let $f:X\to\mathbb{R}$ be a continuous function such that $\mu(f^{-1}(0))<c$ for all $\mu\in\Delta$, and consider the closed set $V:=f^{-1}(0)$. By Lemma 2.3, there is an open set $U\supseteq V$ such that $\mu(U)<c$ for all $\mu\in\Delta$.
Therefore, the set
$\{f\in\mathrm{C}_{\mathbb{R}}(X):\mu(f^{-1}(0))=0,\ \mu\in\Delta\}$
$=$ $\bigcap_{n=1}^{\infty}\{f\in\mathrm{C}_{\mathbb{R}}(X):\mu(f^{-1}(0))<1/n,\ \mu\in\Delta\}$
$=$ $\bigcap_{n=1}^{\infty}Z_{1/n}$
is dense in $\mathrm{C}_{\mathbb{R}}(X)$ (a dense $G_{\delta}$, in fact), and $(X,\Delta)$ has the (SBP).
For the converse, assume that $(X,\Delta)$ has the (SBP), and let $f:X\to\mathbb{R}$ and any $\varepsilon>0$ be given. By Lemma 2.4, there exists a partition of unity $\phi_{i}:X\to[0,1]$, $i=1,2,...,N$, such that
(2.1) $\mu(\bigcup_{i=1}^{N}\phi_{i}^{-1}((0,1))<\varepsilon,\quad\mu\in\Delta,$
and
$\left|f(x)-f(y)\right|<\varepsilon/2,\quad x,y\in\phi_{i}^{-1}(0,1]),\ i=1,...,N.$
For each $i=1,...,N$, pick $x_{i}\in\phi_{i}^{-1}(0,1]$ and a real number $y_{i}$ such that
$y_{i}\neq 0\quad\textrm{and}\quad\left|y_{i}-f(x_{i})\right|<\varepsilon/2.$
Define
$g=\sum_{i=1}^{N}y_{i}\phi_{i}.$
Note that for each $x\in\bigcup_{i=1}^{N}\phi_{i}^{-1}(1)$, one has that $g(x)\neq 0$, and hence
$g^{-1}(0)\subseteq X\setminus\bigcup_{i=1}^{N}\phi_{i}^{-1}(1)=\bigcup_{i=1}^{N}\phi_{i}^{-1}((0,1)).$
By (2.1), direct calculations show that
  1. (1)
    $\left\|f-g\right\|_{\infty}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$,
as desired. ∎
Definition 2.6.
Let $\Delta$ be a closed subset of probability Borel measures on $X$. Then define
$\left\|f\right\|_{2,\Delta}=\sup\{(\int\left|f\right|^{2}d\mu)^{\frac{1}{2}}:\mu\in\Delta\},\quad f\in\mathrm{C}(X).$
Lemma 2.7 (Markov’s inequality).
[…]
where $E_{\varepsilon}=\{x\in X:\left|f(x)\right|>\varepsilon\}$.
Proof.
[…]
Lemma 2.8.
Let $f,g$ be real-valued continuous functions on $X$ satisfying
$\left\|f-g\right\|_{2,\Delta}^{2}<\frac{\varepsilon^{3}}{8}$
and
$\mu(g^{-1}(0))<\frac{\varepsilon}{2},\quad\mu\in\Delta.$
Then, there is a real-valued continuous function $g^{\prime}$ on $X$ such that
  1. (1)
    $\left\|f-g^{\prime}\right\|_{\infty}<\varepsilon$, and
  2. (2)
    $\mu((g^{\prime})^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$.
Proof.
Define
$E_{\varepsilon}=\{x\in X:\left|f(x)-g(x)\right|>\varepsilon/2\}.$
Then, by Markov’s inequality,
$\mu(E_{\varepsilon})\leq\frac{4}{\varepsilon^{2}}\left\|f-g\right\|_{2,\Delta}^{2}<\frac{\varepsilon}{2},\quad\mu\in\Delta.$
Pick an open neighbourhood $U$ of $E_{\varepsilon}^{c}$ such that
$\left|f(x)-g(x)\right|<\varepsilon,\quad x\in U,$
and pick a continuous function $h:X\to[0,1]$ such that
$h|_{E_{\varepsilon}^{c}}=1\quad\mathrm{and}\quad h|_{U^{c}}=0.$
Define
$g^{\prime}(x)=h(x)g(x)+(1-h(x))f(x).$
Then
$\left|f(x)-g^{\prime}(x)\right|<\varepsilon,\quad x\in X.$
Moreover, since
$(g^{\prime})^{-1}(0)\subseteq g^{-1}(0)\cup E_{\varepsilon},$
one has
$\mu((g^{\prime})^{-1}(0))\leq\mu(g^{-1}(0))+\mu(E_{\varepsilon})<\varepsilon,$
as desired. ∎
Using this together with Lemma 2.5, we have the following criterion for the (SBP).
Theorem 2.9.
$(X,\Delta)$ has the (SBP) if, and only if, for any continuous real-valued function $f:X\to\mathbb{R}$ and any $\varepsilon>0$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{2,\Delta}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$; this is equivalent to the existence of $\delta>0$ such that $\tau_{\mu}(\chi_{\delta}(g))<\varepsilon$, $\mu\in\Delta$, where
    $\chi_{\delta}(t)=\left\{\begin{array}[]{ll}1,&|t|<\delta/2,\\ 0,&|t|>\delta,\\ \textrm{linear},&\textrm{otherwise}.\end{array}\right.$
Proof.
By Lemma 2.5 and the fact that $\left\|a\right\|_{2,\Delta}\leq\left\|a\right\|$, we only need to show that if $(X,\Delta)$ has the property that for any continuous real-valued function $f:X\to\mathbb{R}$ and any $\varepsilon>0$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{2,\Delta}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$,
then $(X,\Delta)$ has the (SBP). By Lemma 2.5 again, we only need to show that $(X,\Delta)$ has the property that for any continuous real-valued function $f:X\to\mathbb{R}$ and any $\varepsilon>0$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{\infty}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$.
With the given $\varepsilon>0$, applying the assumption with $f$ and $\min\{\varepsilon^{3}/8,\varepsilon/2\}$, there is a continuous real-valued function $g:X\to\mathbb{R}$ such that
  1. (1)
    $\left\|f-g\right\|_{2,\Delta}<\varepsilon^{3}/8$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon/2$ for all $\mu\in\Delta$.
Then, by Lemma 2.8, there is a real-valued continuous function $g^{\prime}$ on $X$ such that
  1. (1)
    $\left\|f-g^{\prime}\right\|_{\infty}<\varepsilon$, and
  2. (2)
    $\mu((g^{\prime})^{-1}(0))<\varepsilon$ for all $\mu\in\Delta$,
as desired. ∎
We would like to point out the following corollary, which might be interesting on its own.
Corollary 2.10.
Consider $(X,\Delta_{1})$ and $(Y,\Delta_{2})$, and assume there is an embedding $\phi:\mathrm{C}(X)\to\mathrm{C}(Y)$ such that
  1. (1)
    $\Delta_{1}\subseteq\phi^{*}(\Delta_{2})$,
  2. (2)
    $\phi(\mathrm{C}(X))$ is dense in $\mathrm{C}(Y)$ with respect to $\left\|\cdot\right\|_{2,\Delta_{2}}$.
If $(Y,\Delta_{2})$ has the (SBP), then $(X,\Delta_{1})$ also has the (SBP).
Proof.
Identify $\mathrm{C}(X)$ with a sub-C*-algebra of $\mathrm{C}(Y)$, and identify $\Delta_{1}$ with a subset of $\Delta_{2}$.
Let $f:X\to\mathbb{R}$ be continuous, and let $\varepsilon>0$ be arbitrary. Since $(Y,\Delta_{2})$ has the (SBP), by Theorem 2.9, there is $g^{\prime}:Y\to\mathbb{R}$ such that
  1. (1)
    $\left\|\phi(f)-g^{\prime}\right\|_{2,\Delta_{2}}<\varepsilon$,
  2. (2)
    $\mu((g^{\prime})^{-1}(0))<\varepsilon$ for all $\mu\in\Delta_{2}$.
Since $\Delta_{2}$ and $(g^{\prime})^{-1}(0)$ are closed, by Lemma 2.3, there is an open neighbourhood of $(g^{\prime})^{-1}(0)$ with measure uniformly smaller than $\varepsilon$ with respect to $\Delta_{2}$. It then follows that there is $\delta>0$ such that
$\tau(\chi_{\delta}(g^{\prime}))<\varepsilon,\quad\tau\in\Delta_{2}.$
Since $\phi(\mathrm{C}(X))$ is dense inside $\mathrm{C}(Y)$ with respect to $\left\|\cdot\right\|_{2,\Delta_{2}}$, there is a self-adjoint element $g\in\mathrm{C}(X)$ with $\phi(g)$ sufficiently close to $g^{\prime}$ (with respect to $\left\|\cdot\right\|_{2,\Delta_{2}}$) that
$\left\|\phi(f)-\phi(g)\right\|_{2,\Delta_{2}}<\varepsilon\quad\mathrm{and}\quad\tau(\chi_{\delta}(g))<\varepsilon,\quad\mu\in\Delta_{2}.$
In particular, $\mu((g)^{-1}(0))<\varepsilon$ for all $\mu\in\Delta_{1}$. By Theorem 2.9 again, $(X,\Delta_{1})$ has the (SBP). ∎
2.2. Real rank zero
It is useful to characterize the (SBP) using the real rank zero property of a certain sequence algebra. Let us start with the following lemma:
Lemma 2.11.
Assume that $(X,\Delta)$ has the (SBP). Let $f\in\mathrm{C}_{\mathbb{R}}(X)$, and let $\varepsilon>0$. Then there are sequences $(g_{n}),(h_{n})$ in $\mathrm{C}_{\mathbb{R}}(X)$ such that for each $n=1,2,...$,
  1. (1)
    $\left\|f-g_{n}\right\|_{\infty}<\varepsilon$,
  2. (2)
    $\tau_{\mu}(\left|g_{n}h_{n}-1\right|^{2})<1/2^{n}$ for all $\mu\in\Delta$,
  3. (3)
    $\left\|h_{n}\right\|_{\infty}<4/\varepsilon$,
  4. (4)
    $\left\|g_{n}-g_{n+1}\right\|_{2,\Delta}<{1}/{2^{n}}$ and $\left\|h_{n}-h_{n+1}\right\|_{2,\Delta}<{1}/{2^{n}}$.
Proof.
Without loss of generality, one may assume that $\varepsilon<1$ and $\left\|f\right\|=1$.
Choose an open cover $\{U_{1},...,U_{N}\}$ of $X$ such that
$\left|f(x)-f(y)\right|<\varepsilon/2,\quad x,y\in U_{i},\ i=1,...,N.$
By Lemma 2.4, there is a partition of unity $\phi^{(1)}_{i}:X\to[0,1]$, subordinate to $U_{1},U_{2},...,U_{N},$ such that
(2.2) $\mu(\bigcup_{i=1}^{N}(\phi^{(1)}_{i})^{-1}((0,1)))<\varepsilon/2N{({4}/{\varepsilon}+5)^{2}},\quad\mu\in\Delta.$
Consider the collection of open sets
(2.3) $U^{(1)}_{i}=(\phi_{i}^{(1)})^{-1}((0,1]),\quad i=1,2,...,N,$
which is an open cover of $X$ satisfying
$U_{i}^{(1)}\subseteq U_{i},\quad i=1,...,N.$
Using Lemma 2.4 again, one obtains a partition of unity $\phi^{(2)}_{i}:X\to[0,1]$, subordinate to $U^{(1)}_{1},U^{(1)}_{2},...,U^{(1)}_{N}$, such that
$\mu(\bigcup_{i=1}^{N}(\phi^{(2)}_{i})^{-1}((0,1)))<\varepsilon/2^{2}{N({4}/{\varepsilon}+5)^{2}},\quad\mu\in\Delta.$
Noting that, for each $i=1,...,N$, by (2.3), $\phi^{(2)}_{i}(x)=1$ for all $x$ satisfying $\phi^{(1)}_{i}(x)=1$ (so $x\notin U_{j}^{(1)}$, $j\neq i$), one then has
$(\phi_{i}^{(2)})^{-1}((0,1))\subseteq(\phi^{(1)}_{i})^{-1}((0,1))\subseteq\bigcup_{i=1}^{N}(\phi^{(1)}_{i})^{-1}((0,1)),$
and therefore, by (2.2),
$\left\|\phi_{i}^{(1)}-\phi_{i}^{(2)}\right\|_{2,\Delta}<\varepsilon/2N{({4}/{\varepsilon}+5)^{2}}<\varepsilon/2N.$
Repeating this process, one obtains partitions of unity $\phi^{(n)}_{i}:X\to[0,1]$, $i=1,2,...,N$, subordinate to $U_{1},U_{2},...,U_{N}$, such that
(2.4) $\mu(\bigcup_{i=1}^{N}(\phi^{(n)}_{i})^{-1}((0,1)))<\varepsilon/N2^{n}{({4}/{\varepsilon}+5)^{2}},\quad\mu\in\Delta,$
and
(2.5) $\left\|\phi_{i}^{(n)}-\phi_{i}^{(n+1)}\right\|_{2,\Delta}<{\varepsilon}/{2^{n}}N.$
For each $i=1,...,N$, pick $x_{i}\in U_{i}$ and then pick a real number $y_{i}$ such that
$\left|y_{i}\right|>\varepsilon/4\quad\textrm{and}\quad\left|y_{i}-f(x_{i})\right|<\varepsilon/2.$
Define
$g_{n}=\sum_{i=1}^{N}y_{i}\phi^{(n)}_{i}\quad\mathrm{and}\quad h_{n}=\sum_{i=1}^{N}\frac{1}{y_{i}}\phi^{(n)}_{i}.$
Then
$\left\|f-g_{n}\right\|_{\infty}<\varepsilon$
and
$\left|h_{n}(x)\right|\leq\sum_{i=1}^{N}\frac{\phi^{(n)}_{i}(x)}{\left|y_{i}\right|}<\frac{4}{\varepsilon},\quad x\in X.$
Note that $(g_{n}h_{n})(x)=1$ whenever $x\in\bigsqcup_{i=1}^{N}(\phi^{(n)}_{i})^{-1}(1)$, and then, together with (2.4), for all $\mu\in\Delta$,
$\tau_{\mu}(\left|g_{n}h_{n}-1\right|^{2})<(\frac{4\left\|g_{n}\right\|}{\varepsilon}+1)^{2}\cdot\frac{\varepsilon}{2^{n}N(\frac{4}{\varepsilon}+5)^{2}}<(\frac{4(1+\varepsilon)}{\varepsilon}+1)^{2}\frac{1}{2^{n}(\frac{4}{\varepsilon}+5)^{2}}=\frac{1}{2^{n}}.$
By (2.5),
$\left\|g_{n}-g_{n+1}\right\|_{2,\Delta}=\left\|\sum_{i=1}^{N}y_{i}(\phi_{i}^{(n)}-\phi_{i}^{(n+1)})\right\|_{2,\Delta}\leq\sum_{i=1}^{N}\left|y_{i}\right|\left\|\phi_{i}^{(n)}-\phi_{i}^{(n+1)}\right\|_{2,\Delta}\leq 1/2^{n}$
and
$\left\|h_{n}-h_{n+1}\right\|_{2,\Delta}=\left\|\sum_{i=1}^{N}\frac{1}{y_{i}}(\phi_{i}^{(n)}-\phi_{i}^{(n+1)})\right\|_{2,\Delta}\leq\sum_{i=1}^{N}\left|\frac{1}{y_{i}}\right|\left\|\phi_{i}^{(n)}-\phi_{i}^{(n+1)}\right\|_{2,\Delta}\leq 1/2^{n},$
as desired. ∎
Consider $l^{\infty}(\mathrm{C}(X))$ and consider the ideal
$J_{2,\omega,\Delta}:=\{(f_{1},f_{2},...)\in l^{\infty}(\mathrm{C}(X)):\lim_{n\to\omega}(\sup\{\tau(\left|f_{n}\right|^{2}):\tau\in\Delta\})=0\}.$
In general, if $A$ is a C*-algebra and $\Delta$ is a closed set of tracial states of $A$, one also considers $l^{\infty}(A)$ and the ideal
$J_{2,\omega,\Delta}:=\{(a_{1},a_{2},...)\in l^{\infty}(A):\lim_{n\to\omega}(\sup\{\tau(\left|a_{n}\right|^{2}):\tau\in\Delta\})=0\}.$
Theorem 2.12.
The C*-algebra $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ has real rank zero if, and only if, $(X,\Delta)$ has the (SBP).
Proof.
Assume that $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ has real rank zero. Let $f\in\mathrm{C}_{\mathbb{R}}(X)$, and let $\varepsilon>0$ be arbitrary. Consider the image of the constant sequence $\overline{(f,f,...)}\in l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}.$ By the real rank zero assumption, there is an invertible self-adjoint element $\overline{(g_{1},g_{2},...)}\in l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ such that
$\left\|\overline{(f,f,...)}-\overline{(g_{1},g_{2},...)}\right\|_{\infty}<\varepsilon.$
In particular,
(2.6) $\left\|\overline{(f,f,...)}-\overline{(g_{1},g_{2},...)}\right\|_{2,\omega,\Delta}<\varepsilon,$
where
$\left\|(a_{1},a_{2},...)\right\|_{2,\omega,\Delta}=\lim_{n\to\omega}(\sup\{\tau(\left|a_{n}\right|^{2}):\tau\in\Delta\}),\quad(a_{1},a_{2},...)\in l^{\infty}(\mathrm{C}(X)).$
Since $\overline{(g_{1},g_{2},...)}$ is invertible, there is $\delta>0$ such that (with $\chi_{\delta}$ as in Theorem 2.9)
$\chi_{\delta}(\overline{(g_{1},g_{2},...)})=0.$
Hence,
$\mathrm{dist}_{2,\omega,\Delta}((f-g_{1},f-g_{2},...),J_{2,\omega,\Delta})<\varepsilon\quad\mathrm{and}\quad(\chi_{\delta}(g_{1}),\chi_{\delta}(g_{2}),...)\in J_{2,\omega,\Delta},$
and then, with $n$ sufficiently close to $\omega$,
$\left\|f-g_{n}\right\|_{2,\Delta}<\varepsilon\quad\mathrm{and}\quad\tau(\chi_{\delta}(g_{n}))<\varepsilon,\quad\tau\in\Delta.$
By Theorem 2.9, $(X,\Delta)$ has the (SBP).
Now, assume that $(X,\Delta)$ has the (SBP), and let us show that $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ has real rank zero. Let $f\in l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ be a self-adjoint element, and let $\varepsilon>0$ be arbitrary. Pick a self-adjoint representative $(f_{1},f_{2},...)$ of $f$. By Lemma 2.11 (with $f=f_{n}$, $n=1,2,...$), there are self-adjoint elements $g_{1},g_{2},...$ and $h_{1},h_{2},...$ in $\mathrm{C}(X)$ such that for each $n=1,2,...$,
  1. (1)
    $\left\|f_{n}-g_{n}\right\|_{\infty}<\varepsilon$,
  2. (2)
    $\tau_{\mu}(\left|g_{n}h_{n}-1\right|^{2})<1/n$ for all $\mu\in\Delta$, and
  3. (3)
    $\left\|h_{n}\right\|_{\infty}<4/\varepsilon$.
Then, with $g:=\overline{(g_{1},g_{2},...)}$ and $h:=\overline{(h_{1},h_{2},...)}$ in $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$, one has
$\left\|f-g\right\|_{\infty}<\varepsilon\quad\mathrm{and}\quad gh=1.$
That is, $g$ is invertible. Since $\varepsilon$ is arbitrary, this shows that $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ has real rank zero. ∎
[…]
Corollary 2.13.
The pair $(X,\Delta)$ has the (SBP) if (and only if) the invertible self-adjoint elements of $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ are dense in the set of self-adjoint elements with respect to $\left\|\cdot\right\|_{2,\omega,\Delta}$.
Thus, we also have the following (possibly slightly surprising) corollary:
Corollary 2.14.
The C*-algebra $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ has real rank zero if (and only if) its invertible self-adjoint elements are dense in the set of self-adjoint elements with respect to $\left\|\cdot\right\|_{2,\omega,\Delta}$.
The following corollary (directly of Theorem 2.12) is straightforward:
Corollary 2.15.
Consider $(X,\Delta_{1})$ and $(Y,\Delta_{2})$ such that $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta_{1}}\cong l^{\infty}(\mathrm{C}(Y))/J_{2,\omega,\Delta_{2}}$. Then $(X,\Delta_{1})$ has the (SBP) if, and only if, $(Y,\Delta_{2})$ has the (SBP).
Inside $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$, consider the sub-C*-algebra consisting of the classes of bounded sequences of $\mathrm{C}(X)$ which are $\left\|\cdot\right\|_{2,\Delta}$-Cauchy. This C*-algebra is the closure of $\mathrm{C}(X)$ inside $l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ under the norm $\left\|\cdot\right\|_{2,\omega,\Delta}$. Denote it by $\overline{\mathrm{C}(X)}^{\Delta}$ and denote the restriction of $\left\|\cdot\right\|_{2,\omega,\Delta}$ to $\overline{\mathrm{C}(X)}^{\Delta}$ still by $\left\|\cdot\right\|_{2,\Delta}$.
Theorem 2.16.
If the pair $(X,\Delta)$ has the (SBP), then any constant sequence in $\mathrm{C}_{\mathbb{R}}(X)$ can be approximated in $\left\|\cdot\right\|_{\infty}$ by self-adjoint invertible elements of $\overline{\mathrm{C}(X)}^{\Delta}$.
On the other hand, if any constant sequence in $\mathrm{C}_{\mathbb{R}}(X)$ can be approximated in $\left\|\cdot\right\|_{2,\Delta}$ by self-adjoint invertible elements of $\overline{\mathrm{C}(X)}^{\Delta}$, then $(X,\Delta)$ has the (SBP).
Proof.
Assume that any self-adjoint element of $\mathrm{C}(X)\subseteq\overline{\mathrm{C}(X)}^{\Delta}$ is in the $\left\|\cdot\right\|_{2,\omega,\Delta}$ closure of invertible self-adjoint elements. Let $f\in\mathrm{C}_{\mathbb{R}}(X)$, and let $\varepsilon>0$ be arbitrary. Consider the image of the constant sequence $\overline{(f,f,...)}\in l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}.$ By the assumption, there is an invertible self-adjoint element $\overline{(g_{1},g_{2},...)}\in l^{\infty}(\mathrm{C}(X))/J_{2,\omega,\Delta}$ such that
$\left\|\overline{(f,f,...)}-\overline{(g_{1},g_{2},...)}\right\|_{2,\omega,\Delta}<\varepsilon.$
Then there is $\delta>0$ such that
$\chi_{\delta}(\overline{(g_{1},g_{2},...)})=0.$
Hence
$\mathrm{dist}_{2,\omega,\Delta}((f-g_{1},f-g_{2},...),J_{2,\omega,\Delta})<\varepsilon\quad\mathrm{and}\quad(\chi_{\delta}(g_{1}),\chi_{\delta}(g_{2}),...)\in J_{2,\omega,\Delta},$
and so, with sufficiently large $n$, one has
$\left\|f-g_{n}\right\|_{2,\Delta}<\varepsilon\quad\mathrm{and}\quad\tau(\chi_{\delta}(g_{n}))<\varepsilon,\quad\tau\in\Delta.$
By Theorem 2.9, $(X,\Delta)$ has the (SBP).
Now, assume that $(X,\Delta)$ has the (SBP). It follows from Lemma 2.11 that a constant sequence $(f)$ with $f\in\mathrm{C}_{\mathbb{R}}(X)$ can be approximated by an invertible element of $\overline{\mathrm{C}(X)}^{\Delta}$ to within $\varepsilon$, as the sequences $(g_{n})$ and $(f_{n})$ are $\left\|\cdot\right\|_{2,\Delta}$-Cauchy. ∎
Since $\left\|\cdot\right\|_{2,\Delta}$ is dominated by $\left\|\cdot\right\|_{\infty}$, we have the following characterization of the (SBP):
Corollary 2.17.
The pair $(X,\Delta)$ has the (SBP) if, and only if, the invertible self-adjoint elements of $\overline{\mathrm{C}(X)}^{\Delta}$ are dense in the set of self-adjoint elements with respect to $\left\|\cdot\right\|_{2,\omega,\Delta}$.
Remark 2.18.
In contrast to Theorem 2.12, we do not know whether the property (SBP) of $(X,\Delta)$ is characterized by the property (RR0) of $\overline{\mathrm{C}(X)}^{\Delta}$.
3. A stronger version of uniform property $\Gamma$ and the small boundary property
In this section, let us use the characterization of the (SBP) of Theorem 2.9 to connect it to a notion of approximate divisibility on traces (Theorem 3.5).
Recall ([2]) that a C*-algebra $A$ is said to have the uniform McDuff property if for each $n$, there is a unital embedding
$\mathrm{M}_{n}(\mathbb{C})\to(l^{\infty}(A)/J_{2,\omega,\Delta})\cap A^{\prime},$
and the C*-algebra $A$ is said to have uniform property $\Gamma$ if for each $n$, there is a partition of unity
$p_{1},p_{2},...,p_{n}\in(l^{\infty}(A)/J_{2,\omega,\Delta})\cap A^{\prime}$
such that
$\tau(p_{i}ap_{i})=\frac{1}{n}\tau(a),\quad a\in A,\ i=1,2,...,n,\ \tau\in\mathrm{T}(A)_{\omega},$
where $\mathrm{T}(A)_{\omega}$ is the set of traces of $l^{\infty}(A)$ of the form
$\tau((a_{i}))=\lim_{i\to\omega}\tau_{i}(a_{i}),\quad\tau_{i}\in\mathrm{T}(A),$
and $a$ is regarded as the constant sequence $(a)\in l^{\infty}(A)$.
Let us introduce the following stronger versions of these properties, restricting the partitions to belong to a subalgebra $D\subseteq A$:
Definition 3.1.
Let $A$ be a C*-algebra and let $D\subseteq A$ be a sub-C*-algebra. The pair $(D,A)$ is said to have the strong uniform McDuff property if for each $n\in\mathbb{N}$, there is a unital embedding
$\phi:\mathrm{M}_{n}(\mathbb{C})\to(l^{\infty}(A)/J_{2,\omega,\Delta})\cap A^{\prime}$
such that
$\phi(e_{ii})\in l^{\infty}(D)/J_{2,\omega,\Delta},\quad i=1,...,n.$
The pair $(D,A)$ is said to have the strong uniform property $\Gamma$ if for each $n\in\mathbb{N}$, there is a partition of unity
$p_{1},p_{2},...,p_{n}\in(l^{\infty}(D)/J_{2,\omega,\Delta})\cap A^{\prime}$
such that
$\tau(p_{i}ap_{i})=\frac{1}{n}\tau(a),\quad a\in A,\ \tau\in\mathrm{T}(A)_{\omega},$
where $\mathrm{T}(A)_{\omega}$ is as above.
Let us now formulate weaker versions of these two properties, relaxing the approximate centrality of $p_{1},...,p_{n}$, or removing the algebra $A$ altogether.
Definition 3.2.
A pair $(D,A)$, where $A$ is a unital C*-algebra and $D\subseteq A$ is a unital subalgebra, will be said to be (tracially) approximately divisible if there is $K>0$ such that for each $n\in\mathbb{N}$, there is a partition of unity
$p_{1},p_{2},...,p_{n}\in l^{\infty}(D)/J_{2,\omega,\Delta}$
such that
$\tau(p_{i}ap_{i})\leq\frac{1}{n}K\tau(a),\quad a\in D^{+},\ \tau\in\mathrm{T}(A)_{\omega},\ i=1,...,n.$
Similarly, a pair $(D,\Delta)$ (or the pair $(X,\Delta)$), where $D=\mathrm{C}(X)$ and $\Delta$ is a closed set of Borel probability measures, is said to be approximately divisible if there is $K>0$ such that for each $n\in\mathbb{N}$, there is a partition of unity
$p_{1},p_{2},...,p_{n}\in l^{\infty}(D)/J_{2,\omega,\Delta}$
such that
$\tau(p_{i}ap_{i})\leq\frac{1}{n}K\tau(a),\quad a\in D^{+},\ \tau\in\Delta_{\omega},\ i=1,...,n,$
where $\Delta_{\omega}$ is the set of traces of $l^{\infty}(D)$ of the form
$\tau((a_{i}))=\lim_{i\to\omega}\tau_{i}(a_{i}),\quad\tau_{i}\in\Delta.$
Remark 3.3.
It is clear that the strong uniform McDuff property of $(D,A)$ implies the strong uniform property $\Gamma$ (and in particular approximate divisibility).
The property of approximate divisibility can be formulated locally as follows:
Lemma 3.4.
The pair $(D,\Delta)$ is approximately divisible if, and only if, there is $K>0$ such that for any finite set $\mathcal{F}\subseteq D^{+}$, any $\varepsilon>0$, and any $n\in\mathbb{N}$, there are positive contractions $p_{1},...,p_{n}\in D$ such that
  1. (1)
    $\left\|p_{i}p_{j}\right\|_{2,\Delta}<\varepsilon$ if $i\neq j$,
  2. (2)
    $\left\|p_{i}-p_{i}^{2}\right\|_{2,\Delta}<\varepsilon$, $i=1,...,n$,
  3. (3)
    $\left\|1-(p_{1}^{2}+\cdots+p_{n}^{2})\right\|_{2,\Delta}<\varepsilon$,
  4. (4)
    $\tau(p_{i}fp_{i})<\frac{1}{n}K\tau(f)+\varepsilon$, $i=1,...,n$, $f\in\mathcal{F}$, $\tau\in\Delta$.
It turns out that approximate divisibility implies the (SBP):
Theorem 3.5.
Let $D$ be a commutative C*-algebra and let $\Delta$ be a closed set of probability Borel measures. If $(D,\Delta)$ is approximately divisible, then $(D,\Delta)$ has the (SBP).
The converse holds in the case that $\Delta$ is $\mathrm{T}(D\rtimes\Gamma)$ restricted to $D\subseteq D\rtimes\Gamma$, where $\Gamma$ is a discrete amenable group acting freely on $D$.
Let us start with some preparations:
Lemma 3.6.
Let $f\in\mathrm{C}(X)$ be a self-adjoint element, and let $\varepsilon>0$. Then there exist $n\in\mathbb{N}$, and self-adjoint elements $f_{1},f_{2},...,f_{n}\in\mathrm{C}(X)$, such that
  1. (1)
    $\left\|f-f_{i}\right\|_{\infty}<\varepsilon$, $i=1,2,...,n$, and
  2. (2)
    $\frac{1}{n}|\{1\leq i\leq n:f_{i}(x)=0\}|<\varepsilon,\quad x\in X.$
Proof.
Choose an open cover $\mathcal{U}$ of $X$ such that
$\left|f(x)-f(y)\right|<\varepsilon,\quad x,y\in U,\ U\in\mathcal{U}.$
Choose a partition of unity $\{\phi_{U}:U\in\mathcal{U}\}$, subordinate to $\mathcal{U}$, and choose $x_{U}\in U$ for each $U\in\mathcal{U}$. Consider the function
$g:=\sum_{U\in\mathcal{U}}f(x_{U})\phi_{U},$
and it is straightforward to verify that
$\left\|f-g\right\|_{\infty}<\varepsilon.$
It is straightforward to check that $g$ has the factorization
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${N(\mathcal{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$${\mathbb{R}}$
where $N(\mathcal{U})$ is the nerve complex of $\mathcal{U}$ (see, for instance, Definition IX 2.2 of [4] for the definition of the nerve complex of an open cover), the map $\Phi:X\to N(\mathcal{U})$ is given by
$x\mapsto\sum_{U\in\mathcal{U}}\phi_{U}(x)[U],$
and the map $\Psi:N(\mathcal{U})\to\mathbb{R}$ is the linear map
$\sum_{U\in\mathcal{U}}\alpha_{U}[U]\mapsto\sum_{U\in\mathcal{U}}f(x_{U})\alpha_{U}.$
Pick $n>\mathrm{dim}(N(\mathcal{U}))/\varepsilon$, and consider the map
$(\Psi,...,\Psi):N(\mathcal{U})\to\mathbb{R}^{n}.$
By Lemma 5.7 of [6], there is a map
$(\Psi_{1},...\Psi_{n}):N(\mathcal{U})\to\mathbb{R}^{n}$
such that
$\left\|\Psi-\Psi_{i}\right\|_{\infty}<\varepsilon,\quad 1\leq i\leq n$
and
|$|\{1\leq i\leq n:\Psi_{i}(y)=0\}|<\mathrm{dim}(N(\mathcal{U})),\quad y\in N(\mathcal{U}).$
Then
$(f_{1},...,f_{n}):=(\Psi_{1}\circ\Phi,...,\Psi_{n}\circ\Phi)$
satisfies the conclusion of the lemma. ∎
Remark 3.7.
The open cover $\mathcal{U}$ can be chosen to have order $1$, and hence the simplex $N(\mathcal{U})$ to have dimension $1$. Therefore, the number $n$ in the proof can be chosen to be $\lfloor 1/\varepsilon\rfloor+1$, which is independent of $X$.
Corollary 3.8.
Let $f\in\mathrm{C}(X)$ be a self-adjoint element, and let $\varepsilon>0$. Then there exist $n\in\mathbb{N}$, $\delta>0$, and self-adjoint elements $f_{1},f_{2},...,f_{n}\in\mathrm{C}(X)$ such that
  1. (1)
    $\left\|f-f_{i}\right\|_{\infty}<\varepsilon$, $i=1,2,...,n$, and
  2. (2)
    $\frac{1}{n}|\{1\leq i\leq n:|f_{i}(x)|<\delta\}|<\varepsilon,\quad x\in X.$
In particular,
$\frac{1}{n}(\tau((f_{1})_{\delta})+\cdots+\tau((f_{n})_{\delta}))<\varepsilon,\quad\tau\in\mathrm{T}(\mathrm{C}(X)),$
where $(f)_{\delta}=\chi_{\delta}(f)$ with $\chi_{\delta}:\mathbb{R}\to\mathbb{R}$ defined by
$\chi_{\delta}(t)=\left\{\begin{array}[]{ll}1,&|t|<\delta/2,\\ 0,&|t|>\delta,\\ \textrm{linear},&\textrm{otherwise}.\end{array}\right.$
Proof.
By the lemma above, there exist $n\in\mathbb{N}$, and self-adjoint elements $f_{1},f_{2},...,f_{n}\in\mathrm{C}(X)$ such that
  1. (1)
    $\left\|f-f_{i}\right\|_{\infty}<\varepsilon$, $i=1,2,...,n$, and
  2. (2)
    $\frac{1}{n}|\{1\leq i\leq n:|f_{i}(x)|=0\}|<\varepsilon,\quad x\in X.$
Since $f_{1},...f_{n}$ are continuous, for each $x\in X$, there exist a neighbourhood $U_{x}\ni x$ and $\delta_{x}>0$ such that
$\frac{1}{n}|\{1\leq i\leq n:|f_{i}(y)|<\delta_{x}\}|<\varepsilon,\quad y\in U_{x}.$
Then the corollary follows from the compactness of $X$. ∎
It is worth pointing out that the corollary above holds for a general unital C*-algebra.
Corollary 3.9.
Let $A$ be a unital C*-algebra. Let $f\in A$ be a self-adjoint element, and let $\varepsilon>0$. Then there exist $n\in\mathbb{N}$, $\delta>0$, and self-adjoint elements $f_{1},f_{2},...,f_{n}\in A$ such that
  1. (1)
    $\left\|f-f_{i}\right\|_{\infty}<\varepsilon$, $i=1,2,...,n$, and
  2. (2)
    $\frac{1}{n}(\tau((f_{1})_{\delta})+\cdots+\tau((f_{n})_{\delta}))<\varepsilon,\quad\tau\in\mathrm{T}(A).$
Proof.
The statement follows directly from Corollary 3.8 applied to the sub-C*-algebra $D:=\textrm{C*}\{1,f\}$. ∎
Proof of Theorem 3.5.
To show that the pair $(D,\Delta)$ has the (SBP), by Corollary 2.9, it is enough to show that for any self-adjoint element $f\in D$ and any $\varepsilon>0$, there is a self-adjoint element $g\in D$ such that
  1. (1)
    $\|f-g\|_{2,\Delta}<\varepsilon$, and
  2. (2)
    $\mu(g^{-1}(0))<\varepsilon$, $\mu\in\Delta$.
By Corollary 3.8, for the given $\varepsilon$, there exist $n\in\mathbb{N}$ and self-adjoint elements
$f_{1},f_{2},...,f_{n}\in D$
such that
(3.1) $\left\|f-f_{i}\right\|_{\infty}<\varepsilon,\quad i=1,2,...,n,$
and there is $\delta>0$ such that
$\frac{1}{n}(\tau((f_{1})_{\delta})+\cdots+\tau((f_{n})_{\delta}))<\varepsilon/K,\quad\tau\in\mathrm{T}(\mathrm{C}(X)),$
where $K$ is the constant of approximate divisibility. In particular, regarding $(f_{1})_{\delta},...,(f_{n})_{\delta}$ as constant sequences in $D$, we have
(3.2) $\frac{1}{n}(\tau((f_{1})_{\delta})+\cdots+\tau((f_{n})_{\delta}))<\varepsilon/K,\quad\tau\in\Delta_{\omega}.$
By approximate divisibility, there are
$p_{1},p_{2},...,p_{n}\in l^{\infty}(D)/J_{2,\omega,\Delta}$
such that
(3.3) $\tau(p_{i}ap_{i})\leq\frac{1}{n}K\tau(a),\quad a\in D^{+},\ \tau\in\Delta_{\omega}.$
Consider the element
$g:=p_{1}\overline{(f_{1})}p_{1}+\cdots+p_{n}\overline{(f_{n})}p_{n}\in l^{\infty}(D)/J_{2,\omega,\Delta}.$
By (3.1),
(3.4) $\left\|f-g\right\|_{2,\omega,\Delta}=\left\|p_{1}\overline{(f-f_{1})}p_{1}+\cdots+p_{n}\overline{(f-f_{n})}p_{n}\right\|_{2,\omega,\Delta}<\varepsilon.$
Note that, for each $\tau\in\Delta_{\omega}$, by (3.3),
$\tau(p_{i}\overline{((f_{i})_{\delta})}p_{i})\leq\frac{1}{n}K\tau((f_{i})_{\delta}),\quad i=1,...,n,\ \tau\in\Delta_{\omega},$
and hence, together with (3.2),
$\tau((g)_{\delta})$ $=$ $\tau(p_{1}\overline{((f_{1})_{\delta})}p_{1})+\cdots+\tau(p_{n}\overline{((f_{n})_{\delta})}p_{n})$
$\leq$ $\frac{1}{n}K(\tau((f_{1})_{\delta})+\cdots+\tau((f_{n})_{\delta}))$
$<$ $\varepsilon.$
Pick a representative sequence $g=\overline{(g_{k})}$ with $g_{k}$, $k=1,2,...$, self-adjoint elements of $D$. By (3.4) and (3), with some sufficiently large $k$, the function $g_{k}$ satisfies
  1. (1)
    $\left\|f-g_{k}\right\|_{2,\Delta}<\varepsilon$, and
  2. (2)
    $\tau((g_{k})_{\delta})<\varepsilon$, $\tau\in\Delta$,
as desired.
Now, assume that $(D,\mathrm{T}(A))$ has the SBP (Definition 2.2), where $A=D\rtimes\Gamma$ and $(D,\Gamma)$ is free and $\Gamma$ is amenable. It follows from the proof of Theorem 9.4 of [7] that $(D,A)$ has the strong uniform property $\Gamma$, and so $(D,\mathrm{T}(A))$ is approximately divisible. ∎
Corollary 3.10.
Consider a C*-pair $(D,A)$. If $A$ has the strong uniform McDuff property, then $(D,\mathrm{T}(A))$ has the (SBP).
Corollary 3.11.
Consider C*-pairs $(D_{1},A_{1})$ and $(D_{2},A_{2})$, and assume that one of them, say $(D_{1},\mathrm{T}(A_{1}))$, is approximately divisible, then $(D_{1}\otimes D_{2},\mathrm{T}(A_{1}\otimes A_{2}))$ has the (SBP).
Proof.
Note that $\partial\mathrm{T}(A_{1}\otimes A_{2})=\partial\mathrm{T}(A_{1})\times\partial\mathrm{T}(A_{2})$. Then, using Lemma 3.4, it is easy to see that $D_{1}\otimes D_{2}$ is approximately divisible with respect to $\mathrm{T}(A_{1}\otimes A_{2})$ (say, if $(D_{1},A_{1})$ is approximately divisible, then consider the corresponding elements $p_{1}\otimes 1_{D_{2}}$, …, $p_{n}\otimes 1_{D_{2}}$ in $D_{1}\otimes D_{2}$). ∎
Remark 3.12.
Is $(D_{1}\otimes D_{2},A_{1}\otimes A_{2})$ always approximately divisible. (Cf. Theorem 5.6 of [5])
4. The case that $\mathrm{T}(A)$ is a Bauer simplex
As an application of Theorem 3.5, in this section, let us show that if $A=\mathrm{C}(X)\rtimes\mathbb{Z}^{d}$ or $A$ is an AH algebra with diagonal maps, and if $\partial\mathrm{T}(A)$ is compact, then $A$ has the (SBP) (Theorem 4.6), and hence $A$ is $\mathcal{Z}$-stable.
Definition 4.1.
Let $D$ be a unital commutative C*-algebra, and let $\Delta\subseteq\mathrm{T}(D)$ be a closed subset. Let $c=\overline{(c_{1},c_{2},...)}\in l^{\infty}(D)/J_{2,\omega,\Delta}$ be a non-zero positive contraction. Then $\Delta$ is said to be ample with respect to $c$ if for any $\tau_{1},\tau_{2},...\in\Delta$, the limit tracial states
(4.1) $D\ni x\mapsto\frac{1}{(\tau_{n})_{\omega}(c)}(\tau_{n})_{\omega}(xc)=\frac{\lim_{n\to\omega}\tau_{n}(xc_{n})}{\lim_{n\to\omega}\tau_{n}(c_{n})}\in\mathbb{C}$
and
(4.2) $D\ni x\mapsto\frac{1}{1-(\tau_{n})_{\omega}(c)}(\tau_{n})_{\omega}(x(1-c))=\frac{\lim_{n\to\omega}\tau_{n}(x(1-c_{n}))}{1-\lim_{n\to\omega}\tau_{n}(c_{n})}\in\mathbb{C}$
are still in $\Delta$ (when $(\tau_{n})_{\omega}(c)=0$ or $1$, only the trace (4.2) or (4.1), respectively, should be considered), where $(\tau_{n})_{\omega}$ is the limit trace of $(\tau_{n})$ on $l^{\infty}(D)$.
Example 4.2.
Let $A$ be a unital C*-algebra, and let $D\subseteq A$ be a unital commutative sub-C*-algebra. Then the set $\mathrm{T}(A)|_{D}$ is ample with respect to any positive contraction
$c\in(l^{\infty}(D)/J_{2,\omega,\mathrm{T}(A)|_{D}})\cap A^{\prime},$
where $A$ is regarded as the subalgebra of $l^{\infty}(A)/J_{2,\omega,\mathrm{T}(A)}$ consisting of constant sequences, as then, the asymptotic commutativity implies that
$A\ni x\mapsto\frac{1}{(\tau_{n})_{\omega}(c)}(\tau_{n})_{\omega}(xc)=\frac{\lim_{n\to\omega}\tau_{n}(xc_{n})}{\lim_{n\to\omega}\tau_{n}(c_{n})}\in\mathbb{C}$
and
$A\ni x\mapsto\frac{1}{1-(\tau_{n})_{\omega}(c)}(\tau_{n})_{\omega}(x(1-c))=\frac{\lim_{n\to\omega}\tau_{n}(x(1-c_{n}))}{1-\lim_{n\to\omega}\tau_{n}(c_{n})}\in\mathbb{C}$
are tracial states of $A$.
Lemma 4.3.
Consider a pair $(D,\Delta)$ with $\Delta\subseteq\mathrm{T}(D)$ a closed convex subset with $\partial\Delta$ compact. If $\Delta$ is ample with respect to a positive contraction $c=\overline{(c_{n})}\in l^{\infty}(D)/J_{2,\omega,\Delta}$, then
$\lim_{k\to\omega}\sup_{\tau\in\partial\Delta}\left|\tau(ac_{k})-\tau(a)\tau(c_{k})\right|=0,\quad a\in D.$
Proof.
If the statement were not true, then there would exist $a\in D$, $\varepsilon>0$ and a sequence $(\tau_{n})$ in $\Delta$ such that
$\left|\tau_{n}(ac_{n})-\tau_{n}(a)\tau_{n}(c_{n}))\right|\geq\varepsilon,\quad n=1,2,....$
Consider $(\tau_{n})_{\omega}\in\Delta_{\omega}$. Then
(4.3) $(\tau_{n})_{\omega}(ac)\neq(\tau_{n})_{\omega}(a)(\tau_{n})_{\omega}(c),$
and, in particular,
$(\tau_{n})_{\omega}(c)\neq 0,1.$
Note that, for any $x\in D$,
(4.4) $(\tau_{n})_{\omega}(x)=(\tau_{n})_{\omega}(c)\cdot\frac{(\tau_{n})_{\omega}(xc)}{(\tau_{n})_{\omega}(c)}+(1-(\tau_{n})_{\omega}(c))\cdot\frac{(\tau_{n})_{\omega}(x(1-c))}{1-(\tau_{n})_{\omega}(c)}.$
Since $\Delta$ is ample for $c$, we have
$\frac{(\tau_{n})_{\omega}(\ \cdot\ c)}{(\tau_{n})_{\omega}(c)},\quad\frac{(\tau_{n})_{\omega}(\ \cdot\ (1-c))}{1-(\tau_{n})_{\omega}(c)}\in\Delta.$
Since $(\tau_{n})\subseteq\partial\Delta$ and $\partial\Delta$ is compact, $(\tau_{n})_{\omega}|_{D}\in\partial\Delta$; in particular, $(\tau_{n})_{\omega}|_{D}$ is an extreme point of $\Delta$. Therefore (note that $(\tau_{n})_{\omega}(c)\neq 0$), by (4.4),
$(\tau_{n})_{\omega}(x)=\frac{(\tau_{n})_{\omega}(xc)}{(\tau_{n})_{\omega}(c)},\quad x\in D,$
which is
$(\tau_{n})_{\omega}(xc)=(\tau_{n})_{\omega}(x)(\tau_{n})_{\omega}(c),\quad x\in D.$
This contradicts (4.3). ∎
Theorem 4.4.
Consider a pair $(D,\Delta)$ with $\Delta\subseteq\mathrm{T}(D)$ a closed convex subset with $\partial\Delta$ compact. Assume that, for each $n\in\mathbb{N}$, there is a partition of unity $p_{1},p_{2},...,p_{n}\in l^{\infty}(D)/J_{2,\omega,\Delta}$ such that
(4.5) $\tau(p_{i})=\frac{1}{n},\quad\tau\in\Delta_{\omega},\ i=1,...,n,$
where (as before) $\Delta_{\omega}$ is the set of traces of $l^{\infty}(D)$ of the form
$\tau((a_{i}))=\lim_{i\to\omega}\tau_{i}(a_{i}),\quad\tau_{i}\in\Delta,$
and $\Delta$ is ample with respect to each $p_{i}$, $i=1,...,n$. Then, the partition of unity $p_{1},p_{2},...,p_{n}$ above satisfies
(4.6) $\tau(p_{i}ap_{i})=\frac{1}{n}\tau(a),\quad a\in D,\quad\tau\in\Delta_{\omega},\ i=1,...,n.$
So, $(D,\Delta)$ is approximately divisible.
Proof.
Let $a\in D$ and $1\leq i\leq n$ be arbitrary. Pick a representative $(c_{k})\in l^{\infty}(D)$ for $p_{i}$. By Lemma 4.3,
$\lim_{k\to\omega}\sup_{\tau\in\partial\Delta}\left|\tau(ac_{k})-\tau(a)\tau(c_{k})\right|=0,\quad a\in D.$
Together with (4.5), one has
$\lim_{k\to\omega}\sup_{\tau\in\partial\Delta}\left|\tau(ac_{k})-\tau(a)\frac{1}{n}\right|=0,\quad a\in D,$
which implies (4.6). ∎
The theorem above in particular applies to a C*-pair $(D,A)$ where $\Delta=\mathrm{T}(A)|_{D}$. In this case, one actually only has to assume $\mathrm{T}(A)$ to be Bauer (in general this does not obviously imply $\mathrm{T}(A)|_{D}$ is Bauer), and the proof is even shorter:
Theorem 4.5.
Consider a C*-pair $(D,A)$ where $D$ is a unital commutative sub-C*-algebra of $A$, and assume that $\mathrm{T}(A)$ is a Bauer simplex. If $p_{1},p_{2},...,p_{n}\in(l^{\infty}(D)/J_{2,\omega,\mathrm{T}(A)|_{D}})\cap A^{\prime}$, where $n\in\mathbb{N}$, is a partition of unity such that
$\tau(p_{i})=\frac{1}{n},\quad\tau\in\mathrm{T}(A)_{\omega},\ i=1,...,n,$
where $\mathrm{T}(A)_{\omega}$ is the set of traces of $l^{\infty}(A)$ of the form
$\tau((a_{i}))=\lim_{i\to\omega}\tau_{i}(a_{i}),\quad\tau_{i}\in\mathrm{T}(A),$
then,
(4.7) $\tau(p_{i}ap_{i})=\frac{1}{n}\tau(a),\quad a\in A,\quad\tau\in\mathrm{T}(A)_{\omega},\ i=1,...,n.$
Proof.
Let $a\in A$ and $1\leq i\leq n$ be arbitrary. Let $p_{1},p_{2},...,p_{n}\in(l^{\infty}(D)/J_{2,\omega,\mathrm{T}(A)|_{D}})\cap A^{\prime}$ be a partition of unity such that
(4.8) $\tau(p_{i})=\frac{1}{n},\quad\tau\in\mathrm{T}(A)_{\omega},\ i=1,...,n.$
Pick a representative $(b_{k})\in l^{\infty}(D)$ for $p_{i}$. Since $p_{i}\in(l^{\infty}(D)/J_{2,\omega,\mathrm{T}(A)|_{D}})\cap A^{\prime}$, one has
$\lim_{k\to\omega}\left\|ab_{k}-b_{k}a\right\|_{2,\mathrm{T}(A)}=0,\quad a\in A.$
Since $\mathrm{T}(A)$ is a Bauer simplex, by Proposition 3.1 of [2],
$\lim_{k\to\omega}\sup_{\tau\in\partial\mathrm{T}(A)}\left|\tau(ab_{k})-\tau(a)\tau(b_{k})\right|=0,\quad a\in A.$
By (4.8),
$\lim_{k\to\omega}\sup_{\tau\in\mathrm{T}(A)}\left|\tau(b_{k})-\frac{1}{n}\right|=0,$
and hence
$\lim_{k\to\omega}\sup_{\tau\in\partial\mathrm{T}(A)}\left|\tau(ab_{k})-\frac{1}{n}\tau(a)\right|=0,\quad a\in A,$
which implies (4.7). ∎
The uniform Rokhlin property (URP) and the property of comparison of open sets (COS) were introduced in [13]. Every free and minimal dynamical system $(X,\mathbb{Z}^{d})$ has the (URP) and (COS) ([14]); in the case that $\Gamma$ is finitely generated and has subexponential growth, if a free and minimal dynamical system $(X,\Gamma)$ has a Cantor factor, then it has the (URP) and (COS) ([13]). For free and minimal dynamical systems $(X,\Gamma)$ with the (URP) and (COS), if $(X,\Gamma)$ has the (SBP), then the C*-algebra $\mathrm{C}(X)\rtimes\Gamma$ is $\mathcal{Z}$-absorbing ([12]).
Theorem 4.6.
Let $(X,\Gamma)$ be a topological dynamical system with the (URP), where $\Gamma$ is an infinite discrete amenable group. Assume that $\mathcal{M}_{1}(X,\Gamma)$ is a Bauer simplex (but with no restriction on $\mathrm{dim}(\partial\mathcal{M}_{1}(X,\Gamma))$). Then $(X,\Gamma)$ has the (SBP).
In particular, if $(X,\Gamma)$ is free and minimal, and has the (URP) and (COS), the C*-algebra $A:=\mathrm{C}(X)\rtimes\Gamma$ is $\mathcal{Z}$-absorbing, i.e., $A\cong A\otimes\mathcal{Z}$.
Proof.
Since $(X,\Gamma)$ has the property (URP), $\mathcal{M}_{1}(X,\Gamma)$ is canonically isomorphic to $\mathrm{T}(A)$. Therefore $\mathrm{T}(A)$ is a Bauer simplex.
Since $(X,\Gamma)$ has the (URP) and since any discrete amenable group has tilings by Følner sets (Theorem 4.3 of [3]), an argument using Lemma 7.1 of [9] shows that, for any $n\in\mathbb{N}$, there is a partition of unity $p_{1},p_{2},...,p_{n}\in(l^{\infty}(D)/J_{2,\omega,D})\cap A^{\prime}$ such that
$\tau(p_{i})=\frac{1}{n},\quad\tau\in(\mathrm{T}(A))_{\omega},\ i=1,...,n.$
Then, by Theorem 4.4 with $\Delta=\mathrm{T}(A)|_{D}$ (or just Theorem 4.5), one has that $(D,\mathrm{T}(A)|_{D})$ is approximately divisible, and by Theorem 3.5, $(D,\mathrm{T}(A)|_{D})$ has the (SBP).
The second statement then follows from Theorem 5.4 of [11] and Theorem 4.8 of [12]. ∎
Remark 4.7.
If $\left|\Gamma\right|<\infty$ and $(X,\Gamma)$ is minimal, then $X$ must be a finite set, and hence $(X,\Gamma)$ has the (SBP). But the statement does not hold in general without minimality for a finite group $\Gamma$, as the action of the trivial group on $[0,1]$ provides a counterexample.
Since any free and minimal dynamical system $(X,\mathbb{Z}^{d})$ has the (URP) and (COS) (Theorem 4.2 and Theorem 5.5 of [14]), one has the following corollary:
Corollary 4.8.
Let $(X,\mathbb{Z}^{d})$ be a free and minimal topological dynamical system. If $\mathcal{M}_{1}(X,\mathbb{Z}^{d})$ is a Bauer simplex, then $(X,\mathbb{Z}^{d})$ has the (SBP) and the C*-algebra $\mathrm{C}(X)\rtimes\mathbb{Z}^{d}$ is $\mathcal{Z}$-absorbing.
Now, let us consider $A$, an AH algebra with diagonal maps, which is an inductive limit
[…]
where $A_{i}=\bigoplus_{j}\mathrm{M}_{n_{i,j}}(\mathrm{C}(X_{i,j}))$, and the connecting maps have the form
$f\mapsto\mathrm{diag}\{f\circ\lambda_{1},...,f\circ\lambda_{n}\}.$
Since the connecting maps preserve the diagonal subalgebras $D_{i}\subseteq A_{i}$, $i=1,2,...$, their inductive limit $D$ is a commutative subalgebra of $A$. Let us assume that $A$ has no finite-dimensional quotient. Then the matrix ranks $n_{i,j}$ are arbitrarily large as $i\to\infty$, and a standard argument shows that $A$ has the property that, for any $n\in\mathbb{N}$, any finite set $\mathcal{F}\subseteq A$, and any $\varepsilon>0$, there are mutually orthogonal projections $p_{1},...,p_{n}\in D$ such that $p_{1}+\cdots+p_{n}=1$,
$\left|\tau(p_{i})-\frac{1}{n}\right|<\varepsilon,\quad\mathrm{and}\quad\left\|[p_{i},a]\right\|_{2,\tau}<\varepsilon,\quad i=1,...,n,\ a\in\mathcal{F},\ \tau\in\mathrm{T}(A).$
In other words, the conditions of Theorem 4.5 are satisfied for the pair $(D,A)$.
Theorem 4.9.
Let $A$ be a unital separable AH algebra with diagonal maps, and assume that $A$ has no finite-dimensional quotient. If $\mathrm{T}(A)$ is a Bauer simplex, then $(D,\mathrm{T}(A)|_{D})$ has the (SBP). In particular, by Proposition 4.10 below, the C*-algebra $A$ is $\mathcal{Z}$-absorbing if it is simple.
Proof.
Note that (as $A$ is AH) $\mathrm{T}(A)$ is canonically isomorphic to $\mathrm{T}(A)|_{D}$. Since $\mathrm{T}(A)$ is a Bauer simplex, by Theorem 4.4, $(D,\mathrm{T}(A)|_{D})$ is approximately divisible, and so by Theorem 3.5, $(D,\mathrm{T}(A)|_{D})$ has the (SBP). ∎
Just like the transformation group C*-algebra $\mathrm{C}(X)\rtimes\Gamma$ (with the (URP) and (COS)) (see [5] for $\mathbb{Z}$-actions, [13] and [12] for the general case), an AH algebra with diagonal maps which satisfies the (SBP) is also $\mathcal{Z}$-absorbing:
Proposition 4.10.
[…]
$\frac{\mathrm{dim}(Y_{j})}{m_{j}}<\varepsilon,\quad j=1,...,d.$
In particular, $A$ is $\mathcal{Z}$-absorbing.
Proof.
The proof is similar to the proof of Theorem 4.5 of [5]. Let $\varepsilon>0$ be arbitrary, and let $\mathcal{F}\subseteq A$ be a finite subset. Without loss of generality, one may assume that $\mathcal{F}=\mathcal{F}_{0}\cup\mathcal{F}_{1}$ with $\mathcal{F}_{0}\subseteq D\cong\mathrm{C}(X)$ and $\mathcal{F}_{1}$ consisting of finitely many matrix units of a building block. On passing to a finite open cover of $X$ and using the (SBP), the argument of the proof of Lemma 2.4 shows that one may assume further that
$\mathcal{F}_{0}=\{\phi_{1},...,\phi_{L}\},$
where $\phi_{1},...,\phi_{L}$ is a partition of unity of $X$ with the property
$\mu_{\tau}(\phi_{i}^{-1}(0,1))<\frac{\varepsilon}{L},\quad i=1,...,L,\ \tau\in\mathrm{T}(A).$
Then there is a positive contraction $g\in\mathrm{C}(X)$ such that
(4.9) $\tau(g)<\varepsilon\quad\mathrm{and}\quad g\chi_{\varepsilon/4,1-\varepsilon/4}(\phi_{i})=\chi_{\varepsilon/4,1-\varepsilon/4}(\phi_{i}),\quad i=1,...,L,\ \tau\in\mathrm{T}(A),$
where $\chi_{\varepsilon/4,1-\varepsilon/4}$ is a positive function vanishing outside $(\varepsilon/4,1-\varepsilon/4)$, but is $1$ on $[\varepsilon/2,1-\varepsilon/2]$.
With a sufficiently large $k$, there are positive contractions $g^{\prime}\in D_{k}$ and $\phi_{i}\in D_{k}$, $i=1,...,L$, such that
(4.10) $\left\|g-g^{\prime}\right\|_{\infty}<\varepsilon/4\quad\mathrm{and}\quad\left\|\phi^{\prime}_{i}-\phi_{i}\right\|_{\infty}<\varepsilon/4,\quad i=1,...,L.$
Without loss of generality, one may assume that
(4.11) $\tau(g^{\prime})<\varepsilon,\quad\tau\in\mathrm{T}(A_{k}),$
and
$\mathcal{F}_{1}\subseteq A_{k}.$
Then, by (4.9), (4.10), and the fact that $\phi_{i},\phi_{i}^{\prime},g,g^{\prime}$, $i=1,...,L$, commute (they all belong to $D$), one has
(4.12) $\chi_{\varepsilon,1-\varepsilon}((\phi^{\prime}_{i}-\varepsilon)_{+})h_{1-\varepsilon/4}(g^{\prime})=\chi_{\varepsilon,1-\varepsilon}((\phi^{\prime}_{i}-\varepsilon)_{+}),\quad i=1,...,L,$
where $h_{1-\varepsilon/4}$ is the continuous function which is $1$ on $41-\varepsilon/$ 4, 00 at 00, and linear in between. Also note that
(4.13) $\left\|h_{1-\varepsilon/4}(g^{\prime})-g^{\prime}\right\|_{\infty}<\varepsilon/4.$
Then consider
$\theta_{\varepsilon,1-\varepsilon}((\phi_{i}^{\prime}-\varepsilon)_{+}),\quad i=1,...,L,$
where $\theta$ is the positive function which is 00 on $(0,\varepsilon)$, $1$ on $(1-\varepsilon,\infty)$, and linear in between. By (4.12), (4.11), and (4.13),
(4.14) $\mu_{\tau}(\bigcup_{i=1}^{L}(\theta_{\varepsilon,1-\varepsilon}((\phi_{i}^{\prime}-\varepsilon)_{+}))^{-1}(0,1))<\tau(h_{1-\varepsilon/4}(g^{\prime}))<2\varepsilon,\quad\tau\in\mathrm{T}(A_{k}).$
Now, write $A_{k}\cong\bigoplus_{j=1}^{d}\mathrm{M}_{n_{j}}(\mathrm{C}(Z_{j}))$. For each $j=1,...,d$, denote by $V_{j}$ the set of standard matrix units of $\mathrm{M}_{n_{j}}(\mathbb{C})\subseteq A_{k}$. Consider the C*-algebra
$C:=\textrm{C*}\{\theta_{\varepsilon,1-\varepsilon}((\phi_{i}^{\prime}-\varepsilon)_{+})),V_{j}:i=1,...,L,\ j=1,...,d\}\subseteq A_{k}.$
Note that $\mathcal{F}_{0}\subseteq_{3\varepsilon}C$.
Then, by (4.14), an argument similar to the proof of Theorem 4.5 of [5] shows that
$C\cong\bigoplus_{j=1}^{d}\mathrm{M}_{n_{j}}(\mathrm{C}(Z^{\prime}_{j}))$
where $Z_{j}^{\prime}$, $j=1,...,d$ is a metrizable compact space with $\mathrm{dim}(Z^{\prime}_{j})\leq(2\varepsilon)n_{j}$. Since $\mathcal{F}_{0}\subseteq_{3\varepsilon}C$ and $\mathcal{F}_{1}\subseteq A_{k}$, one has that $\mathcal{F}\subseteq_{3\varepsilon}C$. Since $\mathcal{F}$ and $\varepsilon$ are arbitrary, this implies that $A$ locally has slow dimension growth, as desired. ∎